# Towards Real-time Adaptive Anisotropic Image-to-mesh Conversion for Vascular Flow Simulations

### v1

@@ Abstract
Image-to-mesh conversion is handled by a software called CBC3D [27]. CBC3D converts a segmented, multi-labeled image into an isotropic tetrahedral (or mixed element) mesh. The segmented image is discretized using a regular Body-Centered Cubic lattice. The user has the option of adjusting the distance between the vertices of the lattice. Any resulting tetrahedra located outside the object are discarded. Adaptive refinement commences over these high quality elements using red (regular) - green (irregular) subdivisions. A Euclidean Distance Transform is used to determine if an element should be subdivided. CBC3D allows the user to specify an input fidelity parameter which influences how fine the mesh will be near the boundaries. A label-specific fidelity can also be specified to allow for localized refinement in materials of interest (reducing overall element count). Additionally, a mixed element mesh can be created where tetrahedra are merged into hexahedra. The topology of generated surfaces or interfaces between labels is preserved by only merging within homogenous regions where the lattice is uniform. While useful for further reducing element count, the proposed approach only utilizes tetrahedral meshes generated by CBC3D given that CDT3D (the adaptive anisotropic method) only processes tetrahedral background meshes. As seen in our evaluation (see section 5), CDT3D meets the real-time requirement when processing tetrahedral background meshes. The adaptive refinement approach utilized by CBC3D guarantees an output mesh with high quality elements [27]. CBC3D finally deforms the mesh surfaces to their corresponding image boundaries. The final mesh has a smoothness of $C^{0}$ while maintaining the same high quality. Consequently, CBC3D was selected to be the image-to-mesh conversion tool within the pipeline of the proposed approach because it: (1) generates meshes of high quality and fidelity, (2) provides smoothness, and (3) was previously used to discretize patient-specific medical imaging data and STLs of CAD-based stent geometries for use within CFD simulations [40, 26, 25].
3.2 Boundary Layer Generation.
@@
The generation of boundary layers over the high fidelity CBC3D mesh surface is carried out by a software known as AFLR. AFLR is an unstructured mesh generation method used extensively in industry. While it is capable of both isotropic [49] and anisotropic mesh generation [47], it is sequential. Consequently, its use within the proposed approach is only focused on its boundary layer generation capability [48]. AFLR generates boundary layers using a spacing that is normal to the boundary and that grows geometrically. As new points are generated, quality is maintained by checking their distribution (i.e., distance from one another according to a local element length scale). New points are only accepted and inserted into the grid if they produce elements that satisfy dihedral angle and element aspect ratio thresholds. The user has numerous options for controlling the spacing distribution and may also specify the number of layers generated.
3.3 Adaptive Anisotropic Mesh Generation.
@@
The final adaptive anisotropic mesh is generated using a software tool called CDT3D. CDT3D is a multicore cc-NUMA-based mesh generation method that exploits fine-grain parallelism at the cavity level using data decomposition. It is capable of both isotropic [28] and adaptive anisotropic mesh generation [64]. A speculative execution model is implemented, which serves to exploit parallelism “everywhere possible” throughout the mesh generation procedure. Atomic lock instructions are utilized by several of CDT3D’s meshing operations so that different data can be modified concurrently while guaranteeing correctness (i.e., conformity). A lock attempts to acquire the necessary dependencies for a corresponding operation. If unable to do so, any acquired resources are released and the operation is applied on a different set of data. This is repeated until the mesh satisfies spacing and qualitative criteria. There are a number of parallel mesh operations in CDT3D, including: edge collapse, point creation/insertion (which by default uses an advancing front technique for isotropic mesh generation and a centroid-based technique for anisotropic adaptation), CAD projection, local reconnection (edge/face swapping), and point smoothing. See [28] and [64] for more information on their design and implementation.
@@
CDT3D accepts an analytic or discrete metric field as input when performing metric-based anisotropic mesh adaptation. CAD geometries can also be processed and used to perform adaptation (thus allowing the proposed approach to handle both medical I2M conversion and the processing of CAD models for CFD simulations involving stents). CDT3D has been quantitatively and qualitatively evaluated with several benchmarks based on aerospace cases [65, 64]. Not only was it shown to be stable in terms of metric conformity (element shape size and edge length), it also showcased excellent performance when utilizing up to 40 cores on a single multicore node. When coupled with a solver in a CFD simulation pipeline, it captured features of the underlying simulation and occupied only a small fraction of the pipeline’s runtime [64] (a visual example is shown in fig. 2).
4 Method.
@@
The proposed approach involves several steps: (1) utilize CBC3D to convert a segmented image into a high quality and high fidelity background mesh (for anisotropic mesh adaptation), (2) construct an anisotropic metric tensor field, (3) generate a boundary layer grid outside the CBC3D mesh surface using AFLR, and (4) utilize CDT3D to generate an anisotropic mesh volume using the metric tensor field. A visual overview of this pipeline of tools can be seen in fig. 3.
Figure 3: Presented is the pipeline of software tools used to satisfy the medical image-to-mesh conversion requirements for use within a CFD vascular flow simulation. It begins with image-to-mesh conversion (handled by CBC3D). Extracting the medial axis with VMTK [9, 8] is optional, as the method in [21] is only used to build an anisotropic metric tensor field for the purpose of testing this pipeline’s capability to meet the aforementioned requirements (described in section 5.1). A boundary layer grid is generated over the CBC3D mesh using AFLR while the adaptive anisotropic mesh volume is generated by CDT3D (using the CBC3D mesh as a background mesh). The boundary layer grid and anisotropic mesh are merged into one final mesh.
The proposed approach is not bound to any particular method for constructing the metric tensor field. As stated previously, the pipeline is evaluated based on the aforementioned requirements to gauge its suitability for CFD vascular flow simulations. Consequently, our method is tested using two different techniques for constructing a metric tensor field, where each is used with a different anuerysm case. These techniques are described in our experimental setup, section 5.1. The anisotropic metric tensor field and CBC3D background volume mesh (converted from the segmented image) are provided as input to CDT3D, which generates an adaptive anisotropic volume mesh. In this study, surface adaptation is turned off to ensure that the high fidelity surface generated by CBC3D remains intact. The surface is extracted from the CBC3D mesh and is given as input to AFLR to generate a boundary layer grid. AFLR is set to only generate a boundary layer grid, designating the input as an open domain so that only boundary layer elements are generated on the outside of the surface. Surface refinement is also turned off for AFLR (again, for this particular study) to ensure that the high fidelity CBC3D surface remains intact and so that both the boundary layer grid and CDT3D mesh are conforming. The two are merged to create one final output mesh.
5 Results.
5.1 Experimental Setup.
The first aneurysm case is based on a rotational angiography scan of a carotid cavernous aneurysm with image spacing of 1.00 x 1.00 x 1.00 mm3 and an image size of 512 x 512 x 508 voxels3. The CBC3D mesh generated from it can be seen in LABEL:fig:artery1, in addition to its approximate medial axis. The second case is a surface mesh of a middle cerebral artery bifurcation aneurysm given as a Piecewise Linear Complex (PLC) in a .stl format (seen in LABEL:fig:artery2_plc_2). The full pipeline of the proposed approach is tested with the first case while each method attempts to perform mesh generation/adaptation using only the surface mesh for the second case. All codes were compiled using GNU GCC 7.5.0. Data were collected on Old Dominion University’s Wahab cluster using a dual socket node that features two Intel® Xeon® Gold 6148 CPUs @ 2.40 GHz (20 slots) and 384GB of memory. For a performance evaluation, each run in the following experiments was executed five times and the results were averaged.
@@
We evaluate the adaptive anisotropic mesh generation method with metric tensor fields that are constructed using different techniques. For the first case, an anisotropic metric tensor field is constructed based on a method also utilized for non-rigid registration applications intended for image-guided neurosurgery [21]. This method focuses on capturing registration points within an adaptive mesh. A metric tensor field is built by evaluating the k-closest registration points to each mesh vertex using a k-nn search from the VTK library [6]. The Khachiyan algorithm is then used to construct a minimum volume bounding ellipsoid which encompasses the given point set. This ellipsoid has a natural mapping to a 3 x 3 positive definite matrix that can be used as a metric to guide adaptivity. See [21] for more details. To exploit parallelism, an OpenMP construct is utilized when processing the mesh points. For our first test case, we utilize this method using the points of an approximate medial axis of the generated CBC3D mesh (in place of registration points). The approximate medial axis (also called a centerline) is obtained for our first test case using VMTK [9, 8] and can be seen in LABEL:fig:artery1.

### v3

@@ 3.1 Image-to-mesh Conversion
Image-to-mesh conversion in the adaptive anisotropic pipeline is handled by a software called CBC3D [EwCCBC3D]. CBC3D converts a segmented, multi-labeled image into an isotropic tetrahedral (or mixed element) mesh. The segmented image is discretized using a regular Body-Centered Cubic lattice. The user has the option of adjusting the distance between the vertices of the lattice. Any resulting tetrahedra located outside the object are discarded. Adaptive refinement commences over these high quality elements using red (regular) - green (irregular) subdivisions. A Euclidean Distance Transform is used to determine if an element should be subdivided. CBC3D allows the user to specify an input fidelity parameter which influences how fine the mesh will be near the boundaries. A label-specific fidelity can also be specified to allow for localized refinement in materials of interest (reducing overall element count). Additionally, a mixed element mesh can be created where tetrahedra are merged into hexahedra. The topology of generated surfaces or interfaces between labels is preserved by only merging within homogenous regions where the lattice is uniform. While useful for further reducing element count, the proposed approach only utilizes tetrahedral meshes generated by CBC3D given that CDT3D (the adaptive anisotropic method) only processes tetrahedral background meshes. As seen in our evaluation (see section 5), CDT3D meets the real-time requirement when processing tetrahedral background meshes to generate small-size anisotropic meshes (and near real-time when generating large meshes, e.g., 50 million elements). The adaptive refinement approach utilized by CBC3D guarantees an output mesh with high quality elements [EwCCBC3D]. CBC3D finally deforms the mesh surfaces to their corresponding image boundaries. The final mesh has a smoothness of $C^{0}$ while maintaining the same high quality. Consequently, CBC3D was selected to be the image-to-mesh conversion tool within the pipeline of the proposed approach because it: (1) generates meshes of high quality and fidelity, (2) provides smoothness, and (3) was previously used to discretize patient-specific medical imaging data and STLs of CAD-based stent geometries for use within CFD simulations [MelinaNumericalSimulation2016, InVitro2015, HighFidelity2015].
3.2 Boundary Layer Generation
@@
The generation of boundary layers over the high fidelity CBC3D mesh surface is carried out by a software known as AFLR. AFLR is an unstructured mesh generation method used extensively in industry. While it is capable of both isotropic [MarcumAFLR] and anisotropic mesh generation [AnisotropicAFLR], it is sequential. Consequently, its use within the proposed approach is only focused on its boundary layer generation capability [AFLRBoundaryLayer2013]. AFLR generates boundary layers using a spacing that is normal to the boundary and that grows geometrically. As new points are generated, quality is maintained by checking their distribution (i.e., distance from one another according to a local element length scale). New points are only accepted and inserted into the grid if they produce elements that satisfy dihedral angle and element aspect ratio thresholds. The user has numerous options for controlling the spacing distribution and may also specify the number of layers generated.
3.3 Adaptive Anisotropic Mesh Generation
@@
The final adaptive anisotropic mesh is generated using a software tool called CDT3D. CDT3D is a multicore cc-NUMA-based mesh generation method that exploits fine-grain parallelism at the cavity level using data decomposition. It is capable of both isotropic [DrakopoulosCDT3D] and adaptive anisotropic mesh generation [EwCCDT3D]. A speculative execution model is implemented, which serves to exploit parallelism “everywhere possible” throughout the mesh generation procedure. Atomic lock instructions are utilized by several of CDT3D’s meshing operations so that different data can be modified concurrently by threads while guaranteeing correctness (i.e., conformity). A lock attempts to acquire the necessary dependencies for a corresponding operation. If unable to do so, any acquired resources are released and the operation is applied on a different set of data. This is repeated until the mesh satisfies spacing and qualitative criteria. There are a number of parallel mesh operations in CDT3D, including: edge collapse, point creation/insertion (which by default uses an advancing front technique for isotropic mesh generation and a centroid-based technique for anisotropic adaptation), CAD projection, local reconnection (edge/face swapping), and point smoothing. See [DrakopoulosCDT3D] and [EwCCDT3D] for more information on their design and implementation. CDT3D’s adaptation pipeline is organized into two phases - mesh adaptation and quality improvement. A subset of the operations are utilized in each phase [EwCCDT3D]. The mesh adaptation phase is designed to focus on modifying elements so that they conform to point spacing as required by the metric tensor field. The quality improvement phase focuses on improving element shape (i.e., improving element mean ratio). See section 5.2 for details regarding how qualitative criteria are measured in the metric space for our evaluation.
@@
CDT3D accepts an analytic or discrete metric field as input when performing metric-based anisotropic mesh adaptation. CAD geometries can also be processed and used to perform adaptation (thus allowing the proposed approach to handle both medical I2M conversion and potentially the future processing of CAD models for CFD simulations involving stents). CDT3D has been quantitatively and qualitatively evaluated with several benchmarks based on aerospace cases [TsolakisEvaluation, EwCCDT3D]. Not only was it shown to be stable in terms of metric conformity (element shape size and edge length), it also showcased excellent performance when utilizing up to 40 cores on a single multicore node. When coupled with a solver in a CFD simulation pipeline, it captured features of the underlying simulation and occupied only a small fraction of the pipeline’s runtime [EwCCDT3D].
4 Method
We present several contributions to address some of the challenges of medical image-to-mesh conversion.
4.1 I2M Pipeline
@@
The proposed approach involves several steps: (1) utilize CBC3D to convert a segmented image into a high quality and high fidelity background mesh (for anisotropic mesh adaptation), (2) construct an anisotropic metric tensor field, (3) generate a boundary layer grid outside the CBC3D mesh surface using AFLR, and (4) utilize CDT3D to generate an anisotropic mesh volume using the metric tensor field. The boundary layer grid and anisotropic mesh are merged into one final mesh. A visual overview of this pipeline of tools can be seen in Figure 1. A script is utilized, supplying the output of one program as input to the next. Extracting the medial axis with VMTK [VMTK2008, Centerline2003] is optional, as this technique utilized to build an anisotropic metric tensor field only serves as an example for the purpose of testing the pipeline’s robustness.
Figure 1: Presented is the pipeline of software tools used to satisfy the medical image-to-mesh conversion requirements for use within a medical numerical simulation.
@@
The proposed approach is not bound to any particular method for constructing the metric tensor field. As stated previously, the pipeline is evaluated based on the aforementioned requirements to gauge its suitability for medical numerical simulations. Consequently, our method is tested using two different techniques for constructing a metric tensor field, where each is used with a different aneurysm case. For the first case, the metric is constructed based on a method also utilized for non-rigid registration applications intended for image-guided neurosurgery [ComparisonPhysicsBased2023]. This method focuses on capturing registration points within an adaptive mesh. A metric tensor field is built by evaluating the k-closest registration points to each mesh vertex using a k-nearest neighbor search from the VTK library [VTKWebsite]. k is the number of points to find that are closest to a given point. The Khachiyan algorithm [Khachiyan2007] is then used to construct a minimum volume bounding ellipsoid which encompasses the given point set. This ellipsoid has a natural mapping to a 3 x 3 positive definite matrix that can be used as a metric to guide adaptivity. See [ComparisonPhysicsBased2023] for more details. To exploit parallelism, an OpenMP construct is utilized when processing the mesh points. For our first test case, we utilize this method using the points of an approximate medial axis of the generated CBC3D mesh (in place of registration points). The approximate medial axis (also called a centerline) is obtained for our first test case using VMTK [VMTK2008, Centerline2003] and can be seen in Figure LABEL:fig:artery1 of our results.
The anisotropic metric tensor field and CBC3D background volume mesh (converted from the segmented image) are provided as input to CDT3D, which generates an adaptive anisotropic volume mesh. In this study, surface adaptation is turned off to ensure that the high fidelity surface generated by CBC3D remains intact. The surface is extracted from the CBC3D mesh and is given as input to AFLR to generate a boundary layer grid. AFLR is set to only generate a boundary layer grid, designating the input as an open domain so that only boundary layer elements are generated on the outside of the surface (as opposed to within the artery lumen). Such a technique may be useful for fluid-structure interaction (FSI) simulations that couple arterial wall-motion to pulsatile blood flow through the lumen. Mechanical properties of arterial wall tissue (even aneurysmal wall tissue) could be measured and used for the outer layers generated by AFLR. This would allow for calculating metrics such as peak stress within the arterial wall [WallTensionCerebralAneurysms2008, WallsAbdominalAneurysms2025]. While AFLR is used to generate a grid external to the artery lumen for this feasibility study, it can easily be used to generate a grid internal to the lumen for blood flow simulations in future work. The internal AFLR boundary elements can provide increased accuracy of wall shear stress or wall shear stress gradients in the fluid domain. Surface refinement is turned off for AFLR (again, for this particular study) to ensure that the high fidelity CBC3D surface remains intact and so that both the boundary layer grid and CDT3D mesh are conforming. The boundary layer grid (generated outside the artery surface) and the CDT3D mesh (generated within the manifold artery surface) are merged to create one final output mesh. This is done by reading both generated meshes and merging their elements/points to then be output as a single mesh (again, using the VTK library [VTKWebsite]).